# Seek and Solve Reasoning for Table Question Answering

### v1

@@ Abstract
Table-based Question Answering (TQA) involves answering questions based on tabular data. The complexity of table structures and question logic makes this task difficult even for Large Language Models (LLMs). This paper improves TQA performance by leveraging LLMs’ reasoning capabilities. Inspired by how humans solve TQA tasks, we propose a Seek-and-Solve pipeline that instructs the LLM to first seek relevant information and then answer questions. The two stages are integrated at the reasoning level, and their Chain of Thought (CoT) paths are integrated into a coherent Seek-and-Solve CoT (SS-CoT). Furthermore, we present a compact single-stage TQA-solving prompt distilled from the pipeline. Experiments demonstrate that under In-Context Learning settings, using samples with SS-CoT paths as demonstrations, the TQA-solving prompt can effectively guide the LLM to solve complex TQA tasks, resulting in improved performance and reliability. Our results highlight the importance of properly eliciting LLMs’ reasoning capabilities in solving complex TQA tasks.
@@ I Introduction
Some approaches simplify the complex TQA task before letting the model handle it. Common methods include identifying question-relevant data points and using them to simplify the task by highlighting relevant sections of the table [8] or creating simplified sub-tables [9]; see Fig. 2 for an illustration. These operations enhance relevant information while reducing irrelevant ones, making the task easier to handle. Although generally effective, task simplification processes often operate independently of the question-answering mechanism. These methods primarily aim to simplify tasks, frequently overlooking the valuable insights gained during simplification that could help the LLM in answering the question. Consequently, even though the task is simpler, the LLM must reason from scratch. Besides, if relevant information is erroneously omitted during simplification, the LLM may be unable to correct the errors [8]. Both factors suggest that a more holistic approach could further enhance TQA performance.
In another aspect, we revisit the way humans handle complex TQA tasks and find that coherent reasoning is crucial. When dealing with TQA involving complex tables, humans usually follow a two-stage process. First, they analyze the question and understand the table’s semantic structure to seek relevant information. Then, they access the concrete data points and infer the answer step by step. Despite appearing as two separate stages, they are integrated at the level of reasoning, resulting in a coherent reasoning process.
@@
Motivated by these observations, this paper aims to improve TQA performance by leveraging LLMs’ reasoning capabilities. First, we present a novel “Seek-and-Solve” pipeline that directs the LLM’s reasoning in a human-like manner. In the initial “Seek” stage, the LLM is instructed to seek question-relevant information and generate not only the identified relevant information but also the Seek Chain of Thought (CoT) [10] that leads to it. In the “Solve” stage, the Seek-CoT is incorporated into the “Solve” prompt, guiding the LLM to reason consecutively from it rather than starting from scratch to answer the question. Extensive experiments demonstrate that this “Seek-and-Solve” pipeline significantly improves TQA performance and offers higher error tolerance compared to the task simplification process. Second, we present a compact single-stage TQA-solving prompt distilled from the pipeline, leveraging the In-Context Learning (ICL) mechanism to guide the LLM’s reasoning process through demonstrations. The Seek-and-Solve-CoT (SS-CoT), which represents the complete human-like reasoning process for solving a TQA task, can be formed by combining the Seek-CoT and the subsequent Solve-CoT. Experimental results show that when using samples with SS-CoT paths as demonstrations, the TQA-solving prompt can effectively guide the LLM to solve complex TQA tasks. Our findings highlight the importance of properly eliciting LLMs’ reasoning capabilities in solving complex TQA tasks.
@@
The remainder of this paper is organized as follows: Sec. 2 reviews related works. Sec. 3 describes the proposed methods. Sec. 4 presents the experimental results. Finally, Sec. 5 concludes the paper.
@@
Figure 1: Schematic of the Seek-and-Solve pipeline. The ICL prompts are exemplified with one-shot cases.
@@ II Related Works
Prompt Engineering. Prompt engineering is important for eliciting LLM capabilities. CoT prompting [10], which divides reasoning into multi-step formats, has been shown to reduce errors and improve faithfulness. However, not all reasoning paths are equally effective, and some require structured guidance for optimal results [13][14]. ICL [15] is another important technique that provides LLMs with demonstrations to help them solve new tasks. Combining CoT with ICL involves using demonstrations with intermediate reasoning steps to guide the model through complex problem-solving processes using structured examples [16]. In this work, we explore a reasoning path that imitates the human reasoning process and then combine CoT and ICL to instruct an LLM in solving TQA tasks in a dedicated manner.
@@ III Proposed Methods
In this section, we first introduce the tree structure as a tool to model table semantics. Next, we present the “Seek-and-Solve” pipeline. Finally, we present the single-stage TQA-solving prompt.
@@ III-B The Seek-and-Solve pipeline
A schematic of the “Seek-and-Solve” pipeline is shown in Fig. 1. In overview, for a given TQA task, the process involves two LLM calls before arriving at an answer.
Stage-1: Seek. In this stage, we instruct the LLM to enhance its understanding of table semantics and question analysis before performing QA by completing an information seeking task. In specific, we model the table’s row and column header hierarchies as a row tree and a column tree, respectively. The obtained row and column trees are then converted into a list of information, where each item is a tuple linearized from a tree path. Next, we use an ICL prompt (the Stage-1 prompt template in Fig. 1) to instruct the LLM to analyze the question and seek question-relevant tuples (i.e., tree paths) from the list. The resulting Stage-1 response shall be separated into rationale and selected tuples, referred to as Seek-CoT and Seek-Result, respectively, for further use.
@@
Stage-2: Solve. In this stage, we instruct the LLM to answer the question, based on the insights and information gained in Stage 1.
The Seek-CoT is a comprehensive analysis that considers both table semantics and question details. In particular, its intermediate reasoning steps are linked to relevant tree paths. This ensures that each step of the reasoning process is supported by data from the table, potentially improving the overall coherence and accuracy of the question answering. Therefore, we propose allowing the LLM to reason consecutively from the Seek-CoT during the “Solve” stage.
In practice, the Seek-CoT can be used together with the task simplification, creating various effective combinations. We consider two common task simplification ways: providing hints [8] and/or subtables [9] from the Seek-Results. As shown in Fig. 1, the Stage-2 prompt template consists of several slots. By filling different combinations of contents, we can form a corresponding version of the prompt. Specifically, the following variants are considered.
@@
Finally, the answer and the reasoning path, denoted as Solve-Result and Solve-CoT, can be extracted from the Stage-2 response.
@@ III-C The TQA-solving prompt
A compact single-stage TQA-solving prompt can be created by integrating the two prompts of the “Seek-and-Solve” pipeline, as shown in Fig. 3. It takes the full table and the full list of tree paths as inputs and uses an ICL setting to guide the LLM’s reasoning process through demonstrations. In addtion, The Seek-CoT (blue) and its subsequent Solve-CoT (orange) can be integrated to form a Seek-and-Solve-CoT (SS-CoT) that represents a complete reasoning process resembling human reasoning when handling a complex TQA task, as shown in Fig. 3. The SS-CoT distinguishes itself from reasoning paths generated natively by LLMs, referred to as vanilla-CoTs, by using a human-like reasoning process with each step well supported by tree paths. Experiments show that using samples with SS-CoT paths as demonstrations, the TQA-solving prompt can yield performance comparable to “Seek-and-Solve” pipeline while in a single stage.
IV Expriments
@@ IV-B Results with the Seek-and-Solve pipeline
Extensive experiments were performed on HiTab to evaluate the performance of the “Seek-and-Solve” pipeline. As detailed in Sec. III-B, various prompt versions were generated by combining different content elements. The results obtained using Llama-3.1-8B are presented in Tab. I, with each row representing a distinct combination. From the results, several observations can be made. When the Stage-2 LLM reasoned from scratch, simplifying tasks with sub-tables and/or hints yielded significantly better results than those obtained from raw tasks using full tables without any additional information. Conversely, directly providing the full list of tree paths resulted in degraded performance, likely because the LLM struggled to interpret the list without guidance [22].
When the Stage-2 LLM performed reasoning consecutively from Seek-CoT, all prompt versions consistently outperformed their counterparts by a significant margin, demonstrating the value of reasoning-level integration of the two stages. Moreover, the performance gaps between different prompt versions became less significant. This is probably because Seek-CoT incorporates the reasoning behind the task simplifications, making them less necessary.
Furthermore, we analyzed the pipeline’s error tolerance by employing LLMs with varying capacities at its two stages. It is known that a less capable LLM in Stage-1 is more prone to generating errors in Seek-CoT and Seek-Result. Our objective was to determine if a more capable LLM in Stage-2 could mitigate the errors introduced in Stage-1. Tab. II presents the results obtained using Llama-3.1-8B, Mixtral-8x7B, and Llama-3.1-70B. The findings revealed that when certain task simplifications (e.g., subtables and/or hints) were applied, the more capable LLM in Stage-2 struggled to correct the errors from Stage-1, resulting in nontrivial performance gaps. This observation aligned with [8]. Conversely, without task simplification, errors made in Stage-1 were effectively corrected by Llama-3.1-70B in Stage-2, as evidenced by the relatively high performance and small performance gaps, indicating enhanced error tolerance.
Overall, the results presented in Tab. I and Tab. II demonstrated that the “Seek-and-Solve” pipeline led to clear improvements in performance and reliability. This highlighted that the reasoning-level integration of two stages provides significant benefits.
IV-C Results with TQA-solving prompt
@@
We report the performance of the single-stage TQA-solving prompt on HiTab and WikiTQ in Tab. III and Tab. IV, respectively, and compared these results with the state-of-the-art reports.
In HiTab, the TQA-sovling prompt with SS-CoT demonstrations consistently outperformed those with vanilla-CoT demonstrations across all four LLMs, showing significant gains. In WikiTQ, probably due to the relatively flat structure of tables, the results were mixed. While there were performance declines on Mixtral-8x7B and Llama-3.1-8B, the TQA-solving prompt with SS-CoT demonstrations showed improved performance compared to its vanilla-CoT counterparts on Mixtral-8x7B and Llama-3.1-8B. We plan to study this further in the future. Overall, these results indicate that SS-CoT paths properly guided the way LLMs handling complex TQA tasks.
@@ V Conclusion
This paper enhances TQA by leveraging LLM’s reasoning abilities. We proposed a Seek-and-Solve pipeline to guide the LLM to reason in a human-like manner, integrating two stages into a coherent SS-CoT path. Additionally, we presented a single-stage TQA-solving prompt to guide the LLM’s reasoning process through demonstrations with SS-CoT paths. Experimental results show clear improvements in performance and reliability, emphasizing the importance of eliciting reasoning capabilities in LLMs for complex TQA tasks.

### v3

@@ Abstract
The complexities of table structures and question logic make table-based question answering (TQA) tasks challenging for Large Language Models (LLMs), often requiring task simplification before solving. This paper reveals that the reasoning process during task simplification may be more valuable than the simplified tasks themselves and aims to improve TQA performance by leveraging LLMs’ reasoning capabilities. We propose a Seek-and-Solve pipeline that instructs the LLM to first seek relevant information and then answer questions, integrating these two stages at the reasoning level into a coherent Seek-and-Solve Chain of Thought (SS-CoT). Additionally, we distill a single-step TQA-solving prompt from this pipeline, using demonstrations with SS-CoT paths to guide the LLM in solving complex TQA tasks under In-Context Learning settings. Our experiments show that our approaches result in improved performance and reliability while being efficient. Our findings emphasize the importance of eliciting LLMs’ reasoning capabilities to handle complex TQA tasks effectively.
@@ I Introduction
Existing works [6][7][8][9] typically improve complex TQA performance through task simplifications. In our terminology, these multistep approaches usually involve two major stages: the Seek stage and the Solve stage. In the Seek stage, the focus is on simplifying the complex TQA task. Common methods include identifying question-relevant data points and simplifying the task by highlighting relevant sections of the table [8] or creating simplified sub-tables [9]; see Fig. 3 for an illustration. These operations enhance relevant information while reducing irrelevant information, making the task easier to handle. In the subsequent Solve stage, the model solves the simplified tasks instead of the raw tasks to gain benefits.
@@
Although generally effective, the two stages are performed relatively independently. These methods primarily focus on obtaining simplified tasks, often overlooking valuable insights gained during the simplifying process that could aid in subsequent question answering. Consequently, even if the task is simpler, the Solve-stage LLM must reason from scratch. Additionally, if relevant information is erroneously omitted during simplification, the Solve-stage LLM may be unable to correct these errors [8]. Both factors suggest that a more holistic approach could further enhance TQA performance.
Figure 1: An illustration of different pipelines: (a) The pipeline used by existing works, characterized by simplifying the task; (b) The Seek-and-Solve pipeline, which incorporates Seek-CoT into the Solve stage in addition to generating the simplified task; (c) The Seek-and-Solve pipeline, which avoids generating the simplified task by incorporating Seek-CoT into the Solve stage to directly address the raw task; (d) The TQA-solving prompt, a compact single-step prompt.
@@
This paper aims to improve the performance of complex TQA tasks by leveraging the reasoning capabilities of LLMs. Taking a holistic view, we first present a novel Seek-and-Solve pipeline, which integrates the two stages at the reasoning level. In the Seek stage, the LLM is instructed to seek question-relevant information and generate both the simplified task (i.e., Seek-Result) and the Chain of Thought (CoT) [10] that leads to it (i.e., Seek-CoT). In the subsequent Solve stage, the Seek-CoT is incorporated into the Solve prompt, guiding the LLM to reason consecutively from it to solve the raw task. Extensive experiments show that Seek-CoT can be more valuable than Seek-Result. By utilizing Seek-CoT over Seek-Result, the Seek-and-Solve pipeline significantly enhances complex TQA tasks, resulting in improved performance and higher error tolerance.
Secondly, we present a compact single-step TQA-solving prompt distilled from the pipeline, leveraging the In-Context Learning (ICL) mechanism to guide the LLM’s reasoning process through demonstrations. The Seek-and-Solve-CoT (SS-CoT), representing the complete reasoning process for solving a TQA task, is formed by combining the Seek-CoT and the subsequent Solve-CoT. Experimental results show that using samples with SS-CoT paths as demonstrations, this single-step TQA-solving prompt can guide the LLM to achieve performance on par with multistep approaches. Our findings highlight that complex TQA tasks can be addressed both effectively and efficiently once the reasoning capabilities of LLMs are properly elicited.
In summary, rather than boosting TQA performance by simplifying the task, this work explores the potential to enhance TQA performance by eliciting the reasoning capabilities of the LLM. An illustrative comparison is shown in Fig. 1.
@@
The remainder of this paper is organized as follows: Sec. II reviews related works. Sec. III describes the proposed methods. Sec. IV presents the experimental results. Finally, Sec. V concludes the paper.
@@
Figure 2: Schematic of the Seek-and-Solve pipeline. The ICL prompts are exemplified with one-shot cases. Note, the table appears only in Solve-stage prompt.
@@ II Related Works
Prompt Engineering. Prompt engineering is important for eliciting LLM capabilities. CoT prompting [10], which divides reasoning into multi-step formats, has been shown to reduce errors and improve faithfulness. However, not all reasoning paths are equally effective, and some require structured guidance for optimal results [13][14]. ICL [15] is another important technique that provides LLMs with demonstrations to help them solve new tasks. Combining CoT with ICL involves using demonstrations with intermediate reasoning steps to guide the model through complex problem-solving processes using structured examples [16]. In this work, we explore a reasoning path that incorporates domain knowledge and then combine CoT and ICL to instruct an LLM in solving TQA tasks in a dedicated manner.
@@ III Proposed Methods
In this section, we first introduce the tree structure as a tool to model table semantics. Next, we present the Seek-and-Solve pipeline. Finally, we present the compact single-step TQA-solving prompt.
@@ III-B The Seek-and-Solve pipeline
A schematic of the Seek-and-Solve pipeline is shown in Fig. 2. In overview, for a given TQA task, the process involves two LLM calls before arriving at an answer.
The Seek stage. In this stage, we instruct the LLM to enhance its understanding of table semantics and analyze the question before performing QA by completing an information-seeking task. Specifically, we model the table’s row and column header hierarchies as a row tree and a column tree, respectively. The obtained row and column trees are then converted into a list of information, where each item is a tuple linearized from a tree path. Next, we use an ICL prompt (the Seek-stage prompt template in Fig. 2) to instruct the LLM to analyze the question and seek question-relevant tuples (i.e., tree paths) from the list. The resulting Seek-stage response is separated into rationale and selected tuples, referred to as Seek-CoT and Seek-Result, respectively, for further use.
@@
The Solve stage. In this stage, we instruct the LLM to answer the question based on insights and information gained in the Seek stage.
The Seek-CoT is a comprehensive analysis that considers both table semantics and question details. In particular, its intermediate reasoning steps are linked to relevant tree paths. This ensures that each step of the reasoning process is supported by data from the table, potentially improving the overall coherence and accuracy of the question answering. Therefore, we propose allowing the LLM to reason consecutively from the Seek-CoT during the Solve stage.
In practice, the Seek-CoT can be used together with the task simplification, creating various effective combinations. We consider two common task simplification ways: providing hints [8] and/or subtables [9] from the Seek-Results. As shown in Fig. 2, the Solve-stage prompt template consists of several slots. By filling different combinations of contents, we can form a corresponding version of the prompt. Specifically, the following variants are considered.
@@
Finally, the answer and the reasoning path, denoted as Solve-Result and Solve-CoT, can be extracted from the Solve-stage response.
@@ III-C The TQA-solving prompt
A compact single-step TQA-solving prompt can be created by integrating the two prompts of the Seek-and-Solve pipeline, as shown in Fig. 4. It takes the raw task, i.e., the full table and the full list of tree paths, as inputs and uses an In-Context Learning (ICL) setting to guide the LLM’s reasoning process through demonstrations. Additionally, the Seek-CoT (blue) and its subsequent Solve-CoT (orange) can be integrated to form a Seek-and-Solve-CoT (SS-CoT) that represents a complete reasoning process for handling a complex TQA task, as shown in Fig. 4. The SS-CoT distinguishes itself from reasoning paths generated natively by LLMs, referred to as vanilla-CoTs, by grounding each step on relevant table contents.
IV Experiments
@@ IV-A Experimental Settings
IV-B Results with the Seek-and-Solve pipeline
Extensive experiments were conducted on HiTab to evaluate the performance of the Seek-and-Solve pipeline. As detailed in Sec. III-B, various prompt versions were created by combining different content elements. The results obtained using Llama-3.1-8B are presented in Tab. I, with each row representing a distinct combination. Several observations can be made from these results. When the Solve-stage LLM reasoned from scratch, simplifying tasks with Seek-Results (i.e., sub-tables and/or hints) did improve results. However, when the Solve-stage LLM performed reasoning consecutively from Seek-CoT, using Seek-Results became less necessary. This is evidenced by the observation that all prompt versions consistently outperformed their counterparts by a significant margin, and the performance gaps between different prompt versions became less significant. This is likely because Seek-CoT incorporates the reasoning behind the task simplifications, largely covering the value of Seek-Results.
Furthermore, we analyzed the pipeline’s error tolerance by using a more error-prone small LLM at the Seek stage and determining if a more capable Solve-stage LLM could mitigate these errors. The Seek-stage LLM could make two types of errors: one is a Seek-Result Error, caused by omitting relevant information, resulting in a subtable that fails to cover necessary areas or a hint pointing to inaccurate parts. The other is a Seek-CoT Error, which refers to errors in the Seek-stage reasoning process. Tab. II presents the results obtained using Llama-3.1-8B, Mixtral-8x7B, and Llama-3.1-70B. When certain task simplifications based on Seek-Results (i.e., sub-tables and/or hints) were applied, the more capable Solve-stage LLM struggled to correct the errors from the Seek stage, resulting in nontrivial performance gaps. This observation aligned with [8]. Conversely, without task simplification, errors made in the Seek stage were effectively corrected by Llama-3.1-70B in the Solve stage, as evidenced by the relatively high performance and small performance gaps. These findings revealed that the Solve-stage LLM is more effective at correcting Seek-CoT Errors than Seek-Result Errors.
Overall, by using Seek-CoT with raw tasks, the Seek-and-Solve pipeline inherently avoids Seek-Result Errors while eliciting the Solve-stage LLM to solve TQA tasks and correct Seek-CoT Errors, resulting in clear improvements in performance and error tolerance.
@@ IV-B Results with the Seek-and-Solve pipeline
IV-C Results with TQA-solving prompt
We report the performance of the single-step TQA-solving prompt on the HiTab and WikiTQ datasets in Tab. III and Tab. IV, respectively, and compare these results with the state-of-the-art reports. In HiTab, the TQA-solving prompt with SS-CoT demonstrations consistently outperformed those with vanilla-CoT demonstrations across all four LLMs. In WikiTQ, likely due to the relatively flat structure of tables, the results were mixed. While there were performance declines on Mixtral-8x7B and Llama-3.1-8B, the TQA-solving prompt with SS-CoT demonstrations showed improved performance compared to its vanilla-CoT counterparts on Mixtral-8x7B and Llama-3.1-8B. Overall, results show that the TQA-solving prompt properly elicits the reasoning capabilities of LLMs, yielding performance comparable to multistep approaches while operating in a single step, demonstrating both effectiveness and efficiency.
@@ V Conclusion
This work enhances TQA performance by leveraging LLM’s reasoning abilities. We reveal the value of the reasoning process during task simplification and propose a Seek-and-Solve pipeline that integrates two stages at the reasoning level. Additionally, we present a single-step TQA-solving prompt to guide the LLM’s reasoning process through SS-CoT demonstrations. Results demonstrate clear improvements in performance and reliability, underscoring the importance of eliciting reasoning capabilities in LLMs for complex TQA tasks. Broadly, our findings may be inspiring to tasks where reasoning in the prior stage is more valuable than its output.